# GOAL: A Generalist Combinatorial Optimization Agent Learner

### v1

@@ 3.3 Task Representation
To leverage operational similarity between tasks, we need them to have at least representational similarity. Thus, we assume that each instance of size $N$ of a task (CO problem) is defined by a set of $N$ abstract “nodes”, together with features which can be attached at node, edge (node pair) or instance level. Instance features can be replicated and appended to the features of each node, hence can be ignored without loss of generality. Since the construction steps (actions) are returned by the task-specific output adapters (Sec 3.4), there are no restrictions on their format. For simplicity and to cover the problems used in our experiments, we assume in the sequel that each construction step consists of the selection of a node, possibly with an option among a fixed number $K$ of options (dependent on the task: in most cases, $K{=}1$ and the action simply selects a node).
@@
For example, for the path-TSP task, the nodes are the customers to visit plus the origin and destination of the vehicle. There are two proper node features, each holding a binary value indicating whether the node is the origin or the destination, respectively (thus $F_{t}{=}3$ counting node identifier). And there is a single edge feature ($\bar{F}_{t}{=}1$) holding the edge cost (distance, time, energy, etc.). The action is a simple node selection ($K_{t}{=}1$), with origin and destination nodes being masked. In a BQ-MDP transition, the selected node becomes the new origin, and the previous origin is discarded ($N$ is decreased by $1$).
@@ 3.4 Architecture of the Model
The architecture of our model generalizes that of BQ-NCO [16], and is described in Fig. 1 left. It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Additional adapters inside layers can also be introduced, but this paper considers only input-output adapters.. The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features. No positional encoding is used (all relational information is assumed to be captured by $\bar{x}$). The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$. They are self-attention transformer layers [54] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [34] but with a different design.
@@
Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ in self-attention), is an instance of generalized attention [2]. It takes the following inputs: node representations as vectors for $N$ queries, $M$ keys, and $M$ values, with an optional $M{\times}N$ log-binary mask $\mathcal{M}$, as in vanilla attention, plus optional edge representations as $M{\times}N$ “edge” vectors. In each head $h$, each query, key, value vector is first projected, as in vanilla attention, into a common space of dimension $d$, yielding for each node $n{=}1{:}N$ and $m{=}1{:}M$, vectors:
$Q_{n}^{(h)}=Q_{n}\mathbf{W}_{Q}^{(h)},\hskip 14.22636ptK_{m}^{(h)}=K_{m}\mathbf{W}_{K}^{(h)},\hskip 14.22636ptV_{m}^{(h)}=V_{m}\mathbf{W}_{V}^{(h)}.$
The edge vectors are also projected into the same space of dimension $d$, but twice so: once as a key and once as a query, resulting in vectors:
$Q^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{Q},\hskip 14.22636ptK^{\prime(h)}_{mn}=E_{mn}\mathbf{W}^{\prime(h)}_{K}.$
Vanilla attention then builds an $M{\times}N$ score matrix $S_{mn}^{(h)}{=}Q_{n}^{(h)}K_{m}^{(h)\top}$, from which the attention coefficients are computed (using softmax and involving the mask). We simply augment the factors of the scalar product defining the score matrix with the edge projections, and keep the rest unchanged, yielding new representations $r$ of the $N$ nodes:
$\begin{split}r=\sum_{h}\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}V^{(h)}\mathbf{W}_{O}^{(h)\top}\hskip 2.84544pt\\ \text{ with }\hskip 2.84544ptS^{(h)}_{mn}{=}(Q_{n}^{(h)}{+}Q^{\prime(h)}_{mn})(K_{m}^{(h)}{+}K^{\prime(h)}_{mn})^{\top}.\end{split}$ (1)
In GOAL, all the node inputs are of dimension $D$, the edge inputs of dimension $\bar{D}$, and they are all projected into a space of dimension $d$ equal to $D$ divided by the number of heads. Furthermore, no masking is used in the attention blocks.
One well-known limitation of this model is the quadratic complexity bottleneck of attention matrices, which are of size $N^{2}$. In practice though, for many tasks, a simple, task-specific, heuristic can deal with that problem. Essentially, instead of presenting all the $N$ nodes of the input instance to the model, the heuristic selects a priori a subset of the $n$ most promising ones. Thus, the attention matrices never exceed size $n^{2}$, even when $N^{2}$ would be unmanageable. The performance impact is often negligible, and even sometimes improving (maybe due to the dilution of attention coefficient when $N$ is large). For example, with TSP, only the origin, destination and 100-nearest neighbors to the origin are presented.
@@ 3.5 Multi-Type Transformer Architecture
Many tasks (CO problems) involve different types of nodes. For example, scheduling problems often involve “operation” nodes (the things to schedule) and “machine” nodes (which constrain feasible schedules). For GOAL, one solution is to simply treat them all indiscriminately in the backbone. The adapters, being problem specific, are of course free to treat them differently according to their types. This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the edges, though they are often heterogeneous in the input: in the scheduling example, edge features may be precedence constraints between operations, or processing times between machines and operations, and no information at all between machines. Instead of packing all these different types of relations into a single big tensor, with padding values wherever a feature is not defined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges). Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1), operating on different type-compatible subsets of nodes and edges. This is illustrated in Fig. 1 right. Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa. Note that even when edges are not oriented, it is useful to have two cross-attention blocks with transposed matrices as edge embeddings: the projections in (1) are then used to interpret differently the two directions of the edges in each attention block. Finally, observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the possible configurations. This is precisely what enables us to train a unique model to solve various single and multi-type problems.
@@ 4 Experiments
Our experiments seek to answer the following questions: Does the proposed GOAL approach enable learning a single backbone to effectively solve a variety of CO problems? (Sec 4.1); Can we use it as a pretrained model and efficiently fine-tune it to solve new unseen CO problems? (Sec 4.2); and finally, to what extent do the introduced mixed-attention blocks and multi-type architecture contribute to the results? (Sec 4.3). We start by describing the multi-task training of GOAL.
@@
We train GOAL on eight classic and varied CO problems. For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers. Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [19], the Capacitated Vehicle Routing Problem (CVRP) with LKH [19], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [19] the Orienteering Problem (OP) with A4OP [29], the Job Shop Scheduling Problem (JSSP) with ORTools [44], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [24], the Knapsack Problem (KP) with ORTools [44] and the Minimum Vertex Covering Problem (MVC) with FastWVC [1]. The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 10$\times$100 for UMSP (resulting in both cases in 100 operation nodes and 10 machine nodes). For the sake of space, the description of the problems and associated data generation processes is provided in Appendix A.
@@
GOAL is trained by alternating batches from each of the training tasks (round robin), using a single-class cross entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross entropy loss for KP and MVC. Similarly to [16], we create a batch by sampling suffix subproblems for each instance. We use AdamW optimizer [38], with an initial learning rate of 0.0005 and a decay rate every 10th epochs. We trained the model on 8 Nvidia A100 GPU servers, with batch size of 128 for 7 days, allowing the completion of approximately 500 epochs. The best model was selected based on the average performance on a validation set of 512 instances per problem. The model hyperparameters are described in Appendix C.
@@
We consider the following constructive neural models as baselines: AM [30], MDAM [57], POMO [33], Sym-NCO [28], BQ-NCO [16], MVMoE [60], MatNet [34], S2V-DQN [27] and COMPASS [12]. These models are all specialized and trained in single-task mode. While all these models are usually combined with a search procedure, such as beam search [16], simulation-guided beam search  [14], instance augmentation [33], active search [23], for a fair comparison of the underlying neural models, and for sake of space, we only report here the results with a greedy application of each model.
@@ 4.1 Performance on the Training Tasks
[…]
@@
GOAL single-task greedy 0.29% 5s 2.10% 5s 2.54% 5s 0.08% 3s
GOAL multi-task greedy 0.45% 5s 2.85% 5s 3.21% 5s 0.51% 3s
KP100 MVC100 UMSP100x10 JSSP10x10
@@
GOAL single-task greedy 0.10% 3s 0.31% 3s 1.82% 7s 2.65% 6m
GOAL multi-task greedy 0.12% 3s 0.34% 3s 2.28% 7s 3.23% 6m
@@
Generalization to new instance distributions.
Since GOAL is trained with supervised learning on instances of a certain size, it is critical that its performance generalizes to larger instances. Table 2 presents the generalization to instances up to 10 times bigger than the training ones, for a subset of the training tasks. We see an excellent generalization performance on ATSP and CVRP. For the JSSP, the generalization is not as strong but still very good on most distributions. Note that scaling to much larger instances is not a focus of this paper. Recent approaches which do not rely on quadratic attention blocks such as [39] would be more appropriate, although highly specialized to some problems. We believe that the more expensive mixed-attention blocks are crucial for the model’s capacity to handle diverse tasks. We also tested finetuning GOAL to other distribution shifts, namely to the Euclidian distance for TSP and asymmetric distances for CVRP (while GOAL was trained on ATSP and Euclidian CVRP). Results are presented in Appendix B and demonstrate a strong performance.
Train. distribution Generalization

### v2

@@ 3.3 Task Representation
To leverage operational similarity between tasks, we need them to have at least representational similarity. Thus, we assume that each instance of size $N$ of a task (CO problem) is represented by a set of $N$ abstract “nodes”, together with features which can be attached at node, edge (node pair) or instance level. Instance features can be replicated and appended to the features of each node, hence can be ignored without loss of generality. Since the construction steps (actions) are returned by the task-specific output adapters (Sec 3.4), there are no restrictions on their format. For simplicity and to cover the problems used in our experiments, we assume in the sequel that each construction step consists of the selection of a node, possibly with an option among a fixed number $K$ of options (dependent on the task: in most cases, $K{=}1$ and the action simply selects a node).
@@ 3.4 Architecture of the Model
The architecture of our model generalizes that of BQ-NCO [drakulic_bq-nco_2023], and is described in Fig. 1 left. It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Adapters can also be introduced within layers, but this paper considers only input-output adapters.. The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features. No positional encoding is used: all relational information is assumed to be captured by $\bar{x}$. The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$. They are self-attention transformer layers [vaswani_attention_2017] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [kwon_matrix_2022, henderson_transformers_2023] but with a different design.
@@
Our mixed attention block, in the general case where $M$ nodes attend on $N$ nodes ($M{=}N$ only in self-attention), is an instance of generalized attention [andreoli_convolution_2019]. A generalized attention mechanism takes the following inputs: node representations as vectors for $N$ queries, $M$ keys, and $M$ values, with an optional $M{\times}N$ log-binary mask $\mathcal{M}$. In each head $h$, each query, key, value vector is first projected into a common space of dimension $d$, yielding for each $n{=}1{:}N$ and $m{=}1{:}M$, vectors (of dimension $d$):
$Q_{n}^{(h)}$ $=Q_{n}\mathbf{W}_{Q}^{(h)},$ $K_{m}^{(h)}$ $=K_{m}\mathbf{W}_{K}^{(h)},$ $V_{m}^{(h)}$ $=V_{m}\mathbf{W}_{V}^{(h)}.$
Then the key and query vectors are used to compute an $M{\times}N$ score matrix $S$, which is transformed into an attention matrix (by adding the mask and applying softmax) multiplied by the reprojected value vectors to produce the output representation $r$ of the $N$ input query nodes as an $N{\times}D$ matrix:
$r$ $=\sum_{h}\underbrace{\textrm{softmax}_{\textrm{col}}(S^{(h)}{+}\mathcal{M})^{\top}}_{\textrm{attention matrix}}V^{(h)}\mathbf{W}_{O}^{(h)\top}.$
Attention mechanisms then differ in the way they compute the score matrix:
[…]
In vanilla attention, the scores are obtained by simple scalar product of the key and query vectors ((1a)). Our mixed attention mechanism takes as additional input $M{\times}N$ “edge” vectors, projected twice into the common space of dimension $d$, once as key and once as query, resulting in vectors (of dimension $d$):
$Q^{\prime(h)}_{mn}$ $=E_{mn}\mathbf{W}^{\prime(h)}_{Q},$ $K^{\prime(h)}_{mn}$ $=E_{mn}\mathbf{W}^{\prime(h)}_{K}.$ (2)
The score matrix is still obtained as a scalar product of a key and a query vector, but now, each of them is the sum of a node component and an edge component ((1b)). Our mixed attention allows to incorporate arbitrary relational information (edge features) into the core of the attention mechanism. It differs from positional encoding, which attempts to distill a priori the edge information (limited to sequential ordering) into the node representations.
In the GOAL backbone, all the node (resp. edge) inputs to the attention blocks are of dimension $D$ (resp. $\bar{D}$), and they are all projected into a common space of dimension $d$ equal to $D$ divided by the number of heads. No masking is used in the backbone. One well-known limitation of this model is its quadratic complexity, the attention matrices being of size $N^{2}$. In practice though, for many tasks, a simple, task-specific heuristic can deal with that problem by pre-filtering the set of nodes to present to the model (usually much fewer than $N$). Whenever available, we use such heuristics.
Input and output adapters are linear projections, of the task-specific instance features to the backbone’s input embedding (dim. $D$) and from the backbone’s output embedding to task-specific action space. However, when $D$ is large and the number of training tasks is small, there is a risk that the embeddings produced by each task end up occupying complementary subspaces of the embedding space throughout the layers, which would defeat the purpose of effectively sharing parameters across tasks. To prevent this, we simply force the input linear projection to be low rank. Concretely, node (resp. edge) features are first projected into a “small” dimension space, forcing the representations to share dimensions across tasks, then this low dimension representation is plunged into the “large” $D$-dimensional embedding space using a shared codebook (see Fig. 1 left).
@@ 3.5 Multi-Type Transformer Architecture
Many tasks (CO problems) involve different types of nodes. For example, scheduling problems often involve “operation” and “machine” nodes. For GOAL, one solution is to simply treat them all indiscriminately in the backbone. The adapters, being problem specific, are of course free to treat them differently according to their types. This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the nodes and edges, though they are often heterogeneous by nature: in the scheduling example, edge features may carry precedence constraints when between operations, or processing times when between machines and operations, and no information at all when between machines. Instead of packing all these different types of relations into a single vector, with padding values wherever a feature is undefined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges). Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1b), operating on different type-compatible subsets of nodes and edges. This is illustrated in Fig. 1 right. Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa. Observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the configurations: all the attention blocks of a layer, be them self- or cross- attentions, and whatever the types they involve, share the same, layer-specific parameters. This is precisely what allows us to train a unique model to solve arbitrary single- and multi-type problems.
@@ 4 Experiments
Our experiments aim to address the following questions: (i) does the GOAL approach enable learning a single backbone to effectively solve a variety of CO problems (Sec 4.1)? (ii) can it be used as a pretrained model and efficiently fine-tuned to solve new CO problems, unseen at training (Sec 4.2)? (iii) to what extent do our mixed-attention blocks and multi-type architecture contribute to the results (Sec 4.3)? We begin by describing the multi-task training of GOAL.
@@
We train GOAL on eight classic and varied CO problems. For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers. Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem (CVRP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [vidal_hybrid_2022] the Orienteering Problem (OP) with A4OP [kobeaga_efficient_2018], the Knapsack Problem (KP) with ORTools [perron_or-tools_2022], the Minimum Vertex Covering Problem (MVC) with FastWVC [cai_fastwvc_2023], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [huangfu_parallelizing_2018] and the Job Shop Scheduling Problem (JSSP) with ORTools [perron_or-tools_2022]. The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 100$\times$20 for UMSP (resulting in both cases in 100 operation nodes). The description of the problems and the associated data generation processes is provided in Appendix E.
@@
GOAL is trained by alternating batches in random order from each of the training tasks, using a single-class cross-entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross-entropy loss for KP and MVC. Similarly to [drakulic_bq-nco_2023], we create a batch by sampling suffix subproblems for each instance. We use AdamW optimizer [loshchilov_decoupled_2019], with an initial learning rate of 0.0005 and a decay rate of 0.97 every 10th epoch. We trained the model on 8 Nvidia V100 GPU servers, with batch size of 256 for 7 days, completing ca. 400 epochs. The best model was selected based on the average performance on a validation set of 128 instances per problem. The model hyperparameters are described in Appendix A.
@@
We consider the following constructive neural models as baselines: AM [kool_attention_2019], MDAM [xin_multi-decoder_2021], POMO [kwon_pomo_2020], Sym-NCO [kim_sym-nco_2022], BQ-NCO [drakulic_bq-nco_2023], MVMoE [zhou_mvmoe_2024], MatNet [kwon_matrix_2022], S2V-DQN [khalil_learning_2017] and COMPASS [chalumeau_combinatorial_2023]. All models except MVMoE are specialized and trained in single-task mode. They are often combined at test time with various enhancers, such as beam search [drakulic_bq-nco_2023], simulation-guided beam search  [choo_simulation-guided_2022], instance augmentation [kwon_pomo_2020], or active search [hottung_efficient_2022], but for fair comparison, all the results we report are obtained by a pure greedy application of the trained policy at test time.
@@ 4.1 Performance on the Training Tasks
In Table 1 we compare the performance of GOAL to the relevant neural baselines for each problem as well as its single-task version (i.e. GOAL trained separately for each task). First, we observe that the best performance is achieved by the single-task GOAL in 7 out of the 8 tasks, with a computation time comparable to that of the baselines. This proves the versatility and effectiveness of our proposed architecture, even for a specialized model. Second, we see that multi-task GOAL is only slightly worse than the single-task version. These results validate the capability of the GOAL backbone, together with the task-specific adapters, to effectively solve the variety of training tasks. More broadly, they demonstrate that multi-task learning can be a successful paradigm for neural combinatorial optimization.
In addition, in Appendix D, we provide experimental results demonstrating GOAL’s excellent zero-shot generalization on instances up to 10 times larger than the training instances. We can note in particular the strong generalization for ATSP which seems much more challenging for the state-of-the-art baselines. This confirms the effectiveness of our mixed-attention mechanism.
@@
GOAL single-task greedy 0.21% 10s 2.31% 10s 2.79% 10s -0.35% 3s
GOAL multi-task greedy 0.50% 10s 3.06% 10s 2.86% 11s 0.28% 3s
KP100 MVC100 UMSP100x20 JSSP10x10
@@
GOAL single-task greedy 0.10% 3s 0.26% 3s 1.30% 7s 2.70% 15s
GOAL multi-task greedy 0.14% 2s 0.37% 9s 2.01% 7s 3.74% 15s
@@
4.2 Fine-Tuning to New Tasks
In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks. We consider eight diverse new COPs: the Traveling Repairman Problem (TRP), the Prize Collecting TSP (PCTSP), the Open CVRP (OCVRP), the Split Delivery CVRP (SDCVRP), the Sequential Ordering Problem (SOP), the Maximum Covering Location Problem (MCLP), the Maximal Independent Set (MIS) problem and the Open Shop Scheduling Problem (OSSP) (detailed descriptions in Appendix E). GOAL can be fine-tuned to specific problems in two modes. In both, the adapters are trained from scratch, and the backbone is also open to backpropagation (no parameter freeze). The supervised mode is identical to training, except it processes only a few (labeled) instances of the problem for a few steps. Details and results are given in Appendix B. The unsupervised mode is more challenging, but also more realistic, since we may not have access to good-quality solutions for the new problem. We propose to fine-tune the model by iterative imitation-improvement steps between an Apprentice and an Expert, as in ExIt [anthony_thinking_2017]. While Exit relies on a possibly complex tree-search for improvement, we use instead a brute-force but fast method, where the Expert samples multiple solutions (in our case 128) per instance from the Apprentice and selects the best, if it is sufficiently better, and sends it back to the Apprentice (current model) for imitation. This is similar to the self-labeling strategy proposed by corsini_self-labeling_2024 for the JSSP. One could also use more sophisticated experts such as generic local search heuristics or smarter sampling pirnay_self-improvement_2024 to accelerate the finetuning. Performance results are given in Fig 3, which shows that unsupervised fine-tuning systematically outperforms unsupervised training of the whole model from scratch at a comparable processing time budget. We also compare finetuning GOAL to finetuning the single-task models and report the results in Fig 2. We observe that the multi-task model consistently performs as well or better than the “most-related” single-task model (e.g. CVRP model for OCVRP or JSSP model for OSSP), hinting to the synergistic effect of the multi-task training.
Figure 2: Comparison of the supervised fine-tuning of the single- versus multi-task GOAL on four new problems.
Figure 3: Unsupervised fine-tuning of GOAL to eight diverse new tasks. We report the best results over 10 runs both for training from scratch and fine-tuning.
4.3 Ablation Study
Mixed-attention blocks.
We compare our mixed-attention blocks with alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [kwon_matrix_2022] and G2G [henderson_transformers_2023]. More precisely we focus on the ATSP and train three versions of GOAL with the three different variants of mixed-attention. For MatNet, we keep the original hyper-parameters used for solving the ATSP. Fig 4 (left) presents the models performance (optimality gap on a test set of 128 instances) at different points of the training. We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G). Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the larger optimality gap.
Multi-type transformer.
The proposed multi-type transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles. To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node and edge heterogeneity: we simply concatenate the features of the different types of nodes (resp. edges), and use dummy “padding” values whenever a feature is undefined for a given node (resp. edge), so that the original single-type model can be used. We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 (middle) a comparison between the training performance of the single-type and multi-type architectures. Clearly, the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
Sparsification using a codebook.
Fig 4 (right) experimentally confirms the importance, in the fine-tuning phase (supervised mode), of sparsifying the input adapter with a shared codebook. The results are given for the PCTSP, but similar results are available in Fig 6 (Appendix) for the other problems. The shared codebook results in a more stable finetuning (over 10 runs) and to a better optimality gap. While we do not observe a significant effect on the performance on the training tasks, we believe that constraining the rank of the codebook tends to prevent its vectors from specializing for individual tasks seen at training, which would be of limited use when fine-tuning to new tasks.
Figure 4: (left) Performance with different variants of mixed-attention. (middle) Performance of single- vs multi-type GOAL. (right) Fine-tuning of GOAL with and without the codebook.
5 Discussion and Conclusion
In this paper, we propose a generalist neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems. We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on seven CO problems; (ii) a multi-task model, able to solve different problems with the same backbone; and (iii) a pretrained model, which is efficiently fine-tuned to new problems.
The proposed multi-task pre-training relies on imitation learning and therefore the availability of good-quality solutions. While this is not an intrinsic limitation of the GOAL model, we believe that imitation learning is an effective training paradigm for learning multi-task models. This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [sun_smart_2023, reed_generalist_2022]. This is particularly advantageous for CO as we can leverage high-quality solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research. In contrast, for a new potentially more original problem, we show that GOAL can be effectively finetuned without any supervision.
Our approach is limited to problems where solution feasibility is easy to ensure during the construction, which is a common requirement of constructive heuristics. Finally the quadratic complexity of the underlying transformer architecture makes the approach too costly for problem sizes beyond a thousand nodes. Exploring more efficient forms of attention as basis for our mixed-attention blocks would be a promising direction for better scaling.
Appendix A GOAL Hyperparameters
The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [bachlechner_rezero_2021]. The mixed-attention blocks have 8 heads. Low dimensions for nodes and edges are 8 and 4, respectively.
Appendix B Supervised fine-tuning of GOAL
In section 4.2 we describe the process of fine-tuning our multitask model on eight unseen tasks in unsupervised mode. Although fine-tuning without labeled data is useful for transferring to new tasks, it comes at a cost - it is quite expensive. Our experiments show that fine-tuning to the new problems with unsupervised learning takes, on average, a few hours.
On the other hand, we demonstrate that GOAL can be effectively fine-tuned using a very small amount of labeled data, in a short time. Figure 5 presents the performance of fine-tuning GOAL on the same eight problems used in unsupervised tuning. Here, we use only 128 solved (but not necessarily optimal) instances and tune the model with just one training step per instance (more precisely, one training step per tail-subproblem of each instance). In contrast to unsupervised tuning, this method is extremely fast; fine-tuning takes only a few minutes.
Figure 5: Supervised fine-tuning on eight new problems. We report the best results over 10 runs for both training from scratch and fine-tuning.
Appendix C Impact of codebook on fine-tuning
In 4.3 We discussed the importance of sparsifying the input adapter with a shared codebook. Although this feature does not impact training performance, it is very important in the fine-tuning phase. By adding a codebook to the model, fine-tuning becomes more stable and achieves better performance. In the main text, we presented a figure comparing fine-tuning for the PCTSP with models that have and do not have a codebook. Figure 6 illustrates the impact of the codebook on fine-tuning the multitask model across all eight problems we used for transfer. In these experiments, we used light-trained models with and without a codebook. The models are trained on 100K instances, for 200 epochs. For each experiment, we perform 10 inference runs and report the mean with a confidence interval of 100.
Figure 6: Ablation study of codebook
Appendix D Generalization to new instance distributions
Table 2 presents the generalization to instances up to 10 times bigger than the training ones, for a subset of the training tasks. Our approach takes a bit more time than some baselines, but note that scaling to much larger instances is not a focus of this paper. One possible way to avoid the quadratic complexity of attention is by using task-specific heuristics, such as the KNN heuristic for routing problems (proposed by [drakulic_bq-nco_2023], to reduce the size of the input instance and the overall complexity. Another approach could be using methods that do not rely on quadratic attention blocks, such as those proposed by [luo_self-improved_2024], though this method tends to be highly specialized for specific problems.
Our model demonstrates excellent generalization performance in terms of the optimality gap for problems with various edge features, such as ATSP (with an asymmetric distance matrix) and JSSP (with two binary matrices with precedence constraints and job-machine relations).
Train. distr. Generalization